OpenStudy (anonymous):

[Part 1 (4 points)] Show all work in dividing 9x^2-1/8x-4 divided by 3x^2+5x-2/2x^2+1x-1 . [Part 2 (2 points)] Use complete sentences to explain your process.

OpenStudy (campbell_st):

factorise each fraction \[\frac{(3x -1)(3x + 1)}{4(2x -1)} \div \frac{(3x -1)(x +2)}{(2x -1)(x + 1)}\] next flip the 2nd fraction and multiply \[\frac{(3x -1)(3x+1)}{4(2x -1)} \times \frac{(2x -1)(x + 1)}{(3x -1)(x + 2)}\] there are 2 common factors that can be removed (2x -1) and (3x -1) which leaves \[\frac{(3x +1)(x + 1)}{4(x +2)}\] expand for the final solution \[\frac{3x^2 +4x + 1}{4x + 8}\]

OpenStudy (anonymous):

Thanks